A large open tank has two holes in the wall. One is a square hole of side L at a depth Y from the top and other is a circular hole of radius R at a depth 4 y from the top. When the tank is completely filled with water, the quantities of water flowing out per second from both holes are the same. Then R is equal to                                    (a) L/v2p                                                             (c) L
b) 2pL                                                                  (d) L/2p

Grade:12th pass

1 Answers

Aditi Chauhan
9 years ago

Sol:- (a) Equating the rate of flow, we have

v(2gy)*L2 = v(2g * 4y)pR2

[Flow = (area) = (velocity), velocity = v2gx]

Where x =height from top

? L2 = 2pR2 ? R = L/v2p